# Beyond the Dashboard: Investigating Distracted Driver Communication Preferences for ADAS

### v1

@@
Abstract
Distracted driving is a major cause of road fatalities. With improvements in driver (in)attention detection, these distracted situations can be caught early to alert drivers and improve road safety and comfort. However, drivers may have differing preferences for the modes of such communication based on the driving scenario and their current distraction state. To this end, we present a user study (N=147) where videos of simulated driving scenarios were utilized to learn drivers’ preferences for modes of communication and their evolution with the drivers’ changing attention. The survey queried participants preferred modes of communication for scenarios such as collisions or stagnation at a green light. We validate our hypotheses and provide key results that inform the future of communication between drivers and their vehicles. We showcase the different driver preferences based on the nature of the driving scenario and also show that they evolve as the drivers’ distraction state changes. All materials associated with the study can be found at https://sites.google.com/illinois.edu/driver-preference-for-modes.
Index Terms:
@@ Abstract
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
The United States National Highway Traffic Safety Administration reported that 9.8% of road fatalities occurred due to distracted and drowsy drivers in 2021 [1]. This statistic, not including the number of non-fatal cases, could have been reduced through distraction detection. Recent advances have improved driver (in)attention detection by use of methods involving gaze estimation, EEGs, and heart rate monitoring in driver monitoring systems [2, 3, 4, 5]. As distraction detection improves due to these advances, alerting drivers and/or integrating systems for automatic takeovers (i.e. the transfer of control of the vehicle between autonomy and the driver) to prevent such catastrophes becomes paramount in reducing fatalities and other road accidents. Though commercial integration of takeover technologies (e.g. lane keeping assist) has begun with Advanced Driver Assistance Systems (ADAS), drivers find the interventions confusing and uncomfortable without prior explanations [6]. Moreover, most drivers prefer interactive takeover requests over automatic takeovers [7]. Thus, the paradigm for communication for alerts and interactive sessions between drivers and their ADAS is of utmost importance for safe and comfortable driving.
@@
Through improving distraction detection and communication technologies, car manufacturers have introduced safety alerts for drivers who are detected to be distracted or drowsy by the driver monitoring system (e.g. Subaru [8]). As these features are only available in select cars (often expensive), the pool for user feedback that the designers receive is severely limited. These models also use the same modes of communication (i.e. the modality of communicating with driver such as icons and simple beeps) regardless of the drivers abilities to perceive them or their experience using ADAS [8]. This system has led to an increase in over-reliance and decrease in safety. For example, drowsy drivers will often ignore indicators from their cars due to over-reliance [9]. Thus, there is a lack of concrete understanding of driver sentiments and preferences regarding the modes of communication with current solutions.
@@
Therefore, the development and effectiveness of safety alert systems can be significantly improved with a more comprehensive understanding of human-vehicle communication under distraction. While standards for the modes of communication have been established [10, 11], they follow the one mode for all scenarios paradigm. Such a paradigm is impractical as driver psychologies and reaction times have been shown to change under differing (in)attention states [12]. Thus, the questions of if and how driver communication preferences change based on their distraction state is important. Motivated by these questions, we aim to understand the effects of driver distraction on their communication preferences. In particular, we study how driver preferences for the mode of communication changes under different distraction levels and the risk associated with the driving task.
@@
The inferences gained from studying these preferences can then be used to develop an Adaptive Communication Module (AdaCoM) that could alleviate the above problems by autonomously switching the mode of communication based on the driver’s distraction state. Such an idea for human-vehicle interaction (HVI) has existed for decades but has yet to come to fruition due to unreliable perception and the complexities of the human-automata team [13, 14]. The development of these systems becomes more plausible and inevitable as the capabilities in detecting driver distraction improve. While adaptive ADAS such as Adaptive Cruise Control exist, they adapt to factors outside the vehicle and are easily set by the driver. Additionally, without proper user feedback, the framework for driver-ADAS interaction suffers and leads to driver over-reliance and discomfort. Thus, there is a lack of understanding driver preferences for HVI and in-vehicle adaptive communication systems.
@@
Figure 1: Example clips from videos shown in the user survey. Each video is divided into two sections. The top section of the video shows the front view through the windshield from the drivers seat. The bottom section shows the left side view mirror, the rear view mirror, and the right side view mirror from left to right. Video (b) and (c) are not augmented and present all information about the scenario to the participants. Video (a) is augmented to introduce a hyper-fixation distraction.
@@
To address these gaps, we conducted a comprehensive user study [^1] The full survey and videos are linked on the project webpage: https://sites.google.com/illinois.edu/driver-preference-for-modes (N=147) to study drivers’ mode preferences and their evolution. We also study driver sentiments towards AdaCoM systems to motivate their development. We first learn drivers’ preferred modes of communication by way of questionnaires on videos of simulated "real world" driving scenarios. The videos, as shown in Figure 1, were designed to test if drivers preferences depended on the risk associated with the driving scenario and their attention. The driving scenarios depicted in the videos ranged from pedestrian and car collisions to lane deviation and stagnation at green lights. We also asked drivers about their perceived comfort and trust with an adaptive communication module. We analyze the survey responses to show that drivers have different preferences for modes of communication based on the risk involved in the driving scenario. We also verify that drivers’ preferences vary as their distraction state does and that they would want an adaptive communication module in their vehicles.
@@
  1. 1.
    We conduct a user study with 147 participants with diverse experiences to test four hypotheses on how HVI preferences evolve with changing distraction states.
  2. 2.
    We showcase key results and recommendations regarding the use of different modes of communication based on the driving scenario and the drivers distraction state.
  3. 3.
@@
To the best of our knowledge, this is the first user study to directly investigate driver preferences for the modes of communication with ADAS in everyday driving scenarios. This paper is organized as follows. Section II introduces related work. Section III summarizes our hypotheses. We then describe our formal study in Section IV. The results of the user study are presented in Section V. Finally, we conclude the paper and discuss future works and limitations in Section VI.
II Related Work
@@
We provide a brief introduction to the different modes of communication used in human=vehicle communication in this section. We also provide an overview of previous work in adaptive human-automation interaction.
II-A Modalities in Human-Vehicle Communication
@@
The three main modes of communication in Human-Vehicle teams are: (1) Visual; (2) Auditory; and (3) Haptic. Visual displays such as touchscreens and heads-up displays have been used extensively as technology has matured to display non-emergent information [10]. Researchers have focused on the different methods to design these visual interfaces and how their layouts affect drivers’ cognitive load [15, 16, 17]. Conversely, auditory modes such as beeps, ultrasonic bursts, and natural language are recommended for emergent alerts [10, 11, 18]. Haptic modes in different areas of the vehicle such as the steering wheel, seat, and even the gas pedal are also suggested for emergent alerts [19, 20, 21]. Additionally, researchers have aimed to capture the effect of different kinds of auditory signals such as beeps and conversational natural language on drivers cognitive loads [22, 23], and the improved efficiency of non-driving related tasks with tactile modes [24]. Works have also tested the affect of combinations of multiple modes such as speech with visuals [20] and gestures on the steering wheel [25], respectively, and found that the use of the combinations reduced the total cognitive load on the driver.
@@
Even though distracted driving is a major factor affecting road safety, it is often ignored by prior work in human-vehicle communication that instead focuses on the cognitive load placed on the drivers due to different modes as shown above. We address this gap by querying user preferences for the modes of communication under different driver distraction levels. A recent work by Du et al. aims at addressing driver preferences for modes of communication, particularly the different combinations of visual and auditory modes [26]. This work by Du et al. investigates driver preferences under different display modes, types of information, and event criticality. While their work primarily focuses on takeover scenarios, our work is generalized to non-takeover scenarios. Moreover, we also factor in drivers attention level and its affect on their mode preferences. Additionally, we do not restrict ourselves to visual and auditory modes and include other modes such as tactile modes and different types of visual and auditory modes that are omitted from their work. However, unlike Du et al., we do not consider or test the advantages of combining different modes of communication.
II-B Adaptive Human-Automation Interaction
@@
Human-automation adaption is a major field of study at the intersection Psychology and Automotive Sciences. Adaptive systems could benefit human performance if the interaction between the human-automation team was adapted to their task and contextual demands [13]. These adaptive systems would decrease the cognitive load on drivers leading to an improved driving experience [27]. Such adaptive systems would also vastly improve driver engagement as humans tend to lose interest in the task at hand if automation does not adapt quickly [28]. Other works have modeled human-automation interaction with the use of Hidden Markov Models and Markov Decision Processes [29, 30, 31]. These models are trained using simulated data but are yet to be tested on real world systems.
@@
Other works have focused on adapting ADAS to human preferences. Personalizing ADAS has been popular and proposed to improve driving experiences [32, 14]. In particular, Boverie et al.  propose DrivEasy, a framework for adaptive ADAS that behaves similarly to AdaCoM systems [14]. While DrivEasy provides an overview for such an adaptive ADAS framework, no information about the design of adaptive communication is provided. Additionally, driver sentiments towards such a system and its adaptive communication are not highlighted. In our work, we directly address this gap by testing driver comfort and perceived safety with such AdaCoM systems.
@@
Figure 2: An overview of the survey: Participants first fill out a demographics questionnaire, followed by a driving experience questionnaire. Then, participants are randomly assigned one of three tracks that contain distinct videos showcasing the 5 different driving scenarios under different environmental conditions and augmentations. Finally, the survey ends with a questionnaire on their need for and comfort with a system with adaptive communication.
III Hypotheses
In this work, we investigate driver preferences for the mode of communication with the ADAS. Our four hypothesis are:
  1. H1
    Driver preferences for the mode of communication are based on the hazardous nature of the driving scenario. We treat this hypothesis as a confirmation hypothesis as the association of modes of communication with their perception by drivers is standardised by multiple transportation authorities [10, 11].
  2. H2
    Drivers change their preferred mode of communication as their distraction state changes.
  3. H3
    Drivers prefer that their car autonomously adapt its mode of communication to their changing distraction state.
  4. H4
    Vehicles that implement an adaptive policy for its mode of communication are more trusted by drivers.
IV Method
@@
In this section, we describe the procedure and contents of our user study. We first present an overview of the user study. Then, we explain the rationale behind the design of the videos and our survey questionnaires. This study was approved by the institutional review board at the authors’ institute.
IV-A Procedure Overview
The study11footnotemark: 1 was administered online through Google Forms. Participants were recruited through Amazon MTurk, public social media channels such as Reddit and LinkedIn, and email. The use of these channels allowed for a wider variance in participant experiences and backgrounds. Only participants from Amazon MTurk were compensated with \$12 per hour. While the participation of the unpaid respondents was voluntary and due to their interest in the subject matter, we observed no significant differences between the responses of the paid vs unpaid participants.
@@ IV-A Procedure Overview
Participants in the study followed the timeline shown in Figure 2. To begin the study, participants were asked to fill an informed consent form. After submitting the form, they were directed to a demographics questionnaire and a driving experience questionnaire [^2] Details regarding the demographics and driving experience questionnaires are omitted from this paper but can be viewed on the project webpage11footnotemark: 1.. Only participants who were above the age of 18, resided in the United States, and held a drivers license were allowed to proceed further. Participants were then randomly assigned one of three video tracks. Each video track consisted of distinct videos and a questionnaire for the scenario shown in each video. Further details regarding the videos and the questionnaires are presented in Section IV-B. Participants were allowed to view the videos multiple times before answering the questionnaire. Following the video track, participants filled out a questionnaire regarding their opinion on an adaptive communication module for vehicles.
Note that as with similar user studies that were conducted completely online [33, 34], there was no participant oversight by researchers. While responses were scrubbed to exclude faulty responses via completion time analysis (e.g. if participants finished the whole survey very quickly) and response pattern recognition (e.g. if participants provided the same response for all videos), the effect of other factors such as external distractions is not included. After sanitation, 147 user responses were obtained. As a consequence of this nature of the survey, the results presented below must be verified thoroughly through naturalistic studies.
IV-B Videos and Modes of Communication Questionnaire
@@
TABLE I: The detailed video sequences for each track. We show the driving scenario, weather conditions, and if the video was augmented. The slight dominance of scenarios depicting rain in Track 3 was completely incidental.
@@
Scenario Weather Augmented Scenario Weather Augmented Scenario Weather Augmented
@@ 10 Lane Deviation Rain Lane Deviation Night Pedestrian Collision Rain
In order to test H1 and H2, participants were shown videos from a driver’s perspective simulated in the CARLA Driving Simulator [35]. All videos can be viewed on the project webpage11footnotemark: 1. The videos mimic realistic, high-risk scenarios, which could not safely be collected in a real vehicle. As seen in Figure 1, the layout of the videos was designed to show the front view through the windshield, and both side views and the rear view through their respective mirrors.
@@
Each video included one of five driving scenarios at varying levels of risk. The driving scenarios, in order of decreasing risk were: (1) Pedestrian Collisions; (2) Car Collisions; (3) Lane Deviation; (4) Blind Spot Obstruction; (5) Stagnant at a Green Light. The five scenarios were selected due to the ubiquity of ADAS developed to address them (e.g. collision detectors, lane keeping assist systems). The weather in the videos varied between: (a) clear day; (b) rainy day; and (c) night. In total, 15 videos ($3$ weather conditions $\times$ $5$ scenarios) were produced for every combination of driving scenario and weather condition stated above. The number of pedestrians and vehicles and, the environment setting (urban center or highways) were also varied across each video i.e. each video is unique with overlaps in only the type of scenario and the weather condition.
@@
Environmental variations and other agents in the simulation were used to introduce realistic distractions. Particularly, multiple video editing techniques extrapolated from prior work focusing on driver inattention were used to simulate distractions such as: (1) drowsiness; (2) hyper-focus on the road; and (3) noticeable agents in the environment. These distractions were chosen due to their occurrence in driver distraction literature and the ease of implementing them in a simulated environment. Drowsiness was simulated by continually splicing the video with a plain black screen with an increasing duration and frequency to mimic an increase in blinking rate and the change in perception associated with drowsy driving [36]. Inspired by existing literature, we used editing techniques such as zooming in, cropping out, and selectively blurring areas on the videos to simulate hyper-focus on areas in the environment [37]. We also used other dynamic agents in the environment such as children, emergency vehicles, and pedestrians with unique behaviors as sources of distraction outside vehicles. The videos were edited to blur out all views except the aforementioned elements and/or dynamic agents of interest.
@@
Each of the 15 videos were augmented in only one of the styles described above to produce 15 pairs of videos ($3$ weather conditions $\times$ $5$ scenarios $\times$ 2 versions). Therefore, a total of 30 videos were produced and ordered in each of the three tracks. The 15 pairs of videos were divided into three tracks such that each driving scenario appeared exactly once in every track with no overlapping videos across tracks. In particular, each track comprised of 5 videos covering each driving scenario and their augmented versions. All augmented versions of videos appeared before the unedited videos (see Table I). This ordering reduced bias by ensuring that participants did not have a complete view of the scenarios in the augmented versions that they could obtain in the original unedited videos. The videos were also distributed as to have maximum coverage in all tracks for the differing weather conditions and distraction augmentations. Therefore, each scenario was viewed and evaluated independently by every participant.
@@
TABLE II: The Options for the Modes of Communication
@@
After carefully watching each video by imagining themselves as the driver of the vehicle, the participants were asked to answer a mode preference questionnaire. The questionnaire involved ranking the modes of communication listed in Table II and indicating if their vehicle should automatically takeover control. The modes in Table II were selected due to their popular use in vehicles and prior literature [10]. Participants were required to choose only their top three preferred modes for the scenario taking place in every video shown. A Likert scale was used to gather responses for sentiments regarding automatic takeovers for each video.
The questionnaire was designed to obtain the drivers’ mode preferences across "real world" scenarios with differing risk and under distractions. To minimize the effects of individual biases, inferences from the videos were extrapolated using a within-subjects strategy. For this portion of the study, the dependent variables were the ranks associated with each mode and the independent variables were attributes associated with each video i.e. the driving scenario, the weather conditions, and if and how the video was augmented.
We posit that H1 holds true if the there is a change in the number of selections for the modes across the different driving scenarios. Similarly, we posit that H2 holds true if there is a difference between the order of the preferred modes (i.e. rankings of the modes), for the original and augmented versions of the videos per participant. We believe that these tests are robust due to the considerations in the design of the video tracks. Further details regarding the evaluation are discussed in Section V.
IV-C Adaptive Communication Mode Questionnaire

### v2

@@
Abstract.
Distracted driving is a major cause of road fatalities. With improvements in driver (in)attention detection, these distracted situations can be caught early to alert drivers and improve road safety and comfort. However, drivers may have differing preferences for the modes of such communication based on the driving scenario and their current distraction state. To this end, we present an (N=147) where videos of simulated driving scenarios were utilized to learn drivers’ preferences for modes of communication and their evolution with the drivers’ changing attention. The survey queried participants preferred modes of communication for scenarios such as collisions or stagnation at a green light. that inform the future of communication between drivers and their vehicles. We showcase the different driver preferences based on the nature of the driving scenario and also show that they evolve as the drivers’ distraction state changes.
@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; ; isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studies
1. Introduction
@@
The United States National Highway Traffic Safety Administration reported that 9.8% of road fatalities occurred due to distracted and drowsy drivers in 2021 (Stewart, 2023). This statistic, not including the number of non-fatal cases, could have been reduced through distraction detection. Recent advances have improved driver (in)attention detection by use of methods involving gaze estimation, EEGs, and heart rate monitoring in driver monitoring systems (Kashevnik et al., 2021; Shen et al., 2022; Perera et al., 2022; Arakawa, 2021). As distraction detection improves due to these advances, alerting drivers and/or integrating systems for automatic takeovers (i.e. the transfer of control of the vehicle between autonomy and the driver) to prevent such catastrophes becomes paramount in reducing fatalities and other road accidents. Though commercial integration of takeover technologies (e.g. lane keeping assist) has begun with Advanced Driver Assistance Systems (ADAS), drivers find the interventions confusing and uncomfortable without prior explanations (Park et al., 2019). Moreover, most drivers prefer interactive takeover requests over automatic takeovers (Brandenburg and Epple, 2019). Thus, the paradigm for communication for alerts and interactive sessions between drivers and their ADAS is of utmost importance for safe and comfortable driving.
@@
Through improving distraction detection and communication technologies, car manufacturers have introduced safety alerts for drivers who are detected to be distracted or drowsy by the driver monitoring system (e.g. Subaru (Corporation, 2020)). As these features are only available in select (often expensive) cars, the pool for user feedback that the designers receive is severely limited. These models also use the same modes of communication (i.e. the modality of communicating with driver such as icons and simple beeps) regardless of the driver’s abilities to perceive them or their experience using ADAS (Corporation, 2020; DeGuzman and Donmez, 2021). This system has led to an increase in over-reliance and decrease in safety. For example, drowsy drivers will often ignore indicators from their cars due to over-reliance (Dunn et al., 2019). Thus, there is a lack of concrete understanding of driver sentiments and preferences regarding the modes of communication with current solutions.
@@
Additionally, the development and effectiveness of safety alert systems can be significantly improved with a more comprehensive understanding of human-vehicle communication under distraction. While standards for the modes of communication have been established (Campbell et al., 2018; Lerner et al., 2015), they follow the one mode for all scenarios paradigm. Such a paradigm is impractical as driver psychologies and reaction times have been shown to change under differing (in)attention states (Stutts et al., 2003). Thus, the questions of if and how driver communication preferences change based on their distraction state is important.
@@
the inferences gained from studying these preferences can then be used to an Adaptive Communication Module (AdaCoM) that could alleviate the above problems by autonomously switching the mode of communication based on the driver’s distraction state. Such an idea for human-vehicle interaction (HVI) has existed for decades but has yet to come to fruition due to unreliable perception and the complexities of the human-automation team (Hancock et al., 2013; Boverie et al., 2011). The development of these systems becomes more plausible and inevitable as the capabilities in detecting driver distraction improve. While adaptive ADAS such as Adaptive Cruise Control exist, they adapt to factors outside the vehicle and are easily set by the driver.
(a) A clear day
(b) A rainy day
(c) Night time driving
@@
Figure 1. Example clips from videos shown in the user survey. Each video is divided into two sections. The top section of the video shows the front view through the windshield from the driver’s seat. The bottom section shows the left side view mirror, the rear-view mirror, and the right side view mirror from left to right. Video (b) and (c) are not augmented and present all information about the scenario to the participants. Video (a) is augmented to introduce a hyper-fixation distraction.
[…]
[Three clips from videos in the survey showing different driving scenarios in varying weather conditions and environments.]Three clips from videos showing a driving on a highway on a rainy day with traffic, driving on a highway on a clear day, and driving at night with a pedestrian crossing the road from left to right respectively.
 [^1] The full survey and videos are linked on the project webpage: https://sites.google.com/view/beyond-the-dashboard/
@@
(N=147) and study drivers’ mode preferences and their evolution. We also study driver sentiments towards AdaCoM systems to further motivate their development. We first learn drivers’ preferred modes of communication by way of questionnaires on videos of simulated real-world driving scenarios were designed to directly test if drivers’ preferences depended on the risk associated with the driving scenario and their (in)attention. The driving scenarios depicted in the videos ranged from pedestrian and car collisions to lane deviation and stagnation at green lights. We also asked drivers about their perceived comfort with and trust in an adaptive communication module. We analyze the survey responses to show that drivers have different preferences for modes of communication based on the risk involved in the driving scenario. We also verify that drivers’ preferences vary as their distraction state does and that they would want an adaptive communication module in their vehicles.
@@
  1. (1)
    We conduct a user with 147 participants with diverse experiences to on how HVI preferences evolve with changing distraction states.
  2. (2)
    We showcase key results and recommendations regarding the use of different modes of communication based on the driving scenario and the driver’s distraction state.
  3. (3)
@@
To the best of our knowledge, this is the first user to directly investigate driver preferences for the modes of communication with ADAS in everyday driving scenarios. This paper is organized as follows: Section 2 introduces related work. We then describe our in Section 3. The results of the study are presented in Section 4. Finally, we conclude with a discussion on future works and limitations in Section 5.
2. Related Work
@@
We provide a brief introduction to the different modes of communication used in human-vehicle communication in this section. We also provide an overview of previous work in adaptive human-automation interaction as well as user study methodologies in automotive research.
2.1. Modalities in Human-Vehicle Communication
@@
The three main modes of communication in Human-Vehicle teams are: (1) Visual; (2) Auditory; and (3) Haptic. Visual displays such as touchscreens and heads-up displays have been used extensively as technology has matured to display non-emergent information (Campbell et al., 2018). Researchers have focused on the different methods to design these visual interfaces and how their layouts affect drivers’ cognitive load (Olaverri-Monreal et al., 2014; Berger et al., 2022; Topliss et al., 2019). Conversely, auditory modes such as beeps, ultrasonic bursts, and natural language are recommended for emergent alerts (Campbell et al., 2018; Lerner et al., 2015; Harrington et al., 2018). Haptic modes in different areas of the vehicle such as the steering wheel, seat, and even the gas pedal are also suggested for emergent alerts (Meck and Precht, 2021; Kim and Ji, 2019; de Ruiter and Alonso, 2019). Additionally, researchers have aimed to capture the effect of different kinds of auditory signals such as beeps and natural language on drivers’ cognitive loads (Wang et al., 2022; Wong et al., 2019), and the improved efficiency of non-driving related tasks with tactile modes (Vo and Brewster, 2020). Works have also tested the effect of combinations of multiple modes and found that the use of the combinations reduced the total cognitive load on drivers (e.g. speech with visuals (Kim and Ji, 2019), gestures on the steering wheel (Cui et al., 2021)).
@@
Even though distracted driving is a major factor affecting road safety, it is often ignored by prior work in human-vehicle communication that instead focuses on the cognitive load placed on the drivers due to different modes as shown above. We address this gap by querying user preferences for the modes of communication under different driver distraction levels. A recent work by  Du et al. (2021) aims at addressing driver preferences for modes of communication, particularly the different combinations of visual and auditory modes. This work by Du et al. (2021) investigates driver preferences under different display modes, types of information, and event criticality. While their work primarily focuses on takeover scenarios, our work is generalized to non-takeover scenarios. Moreover, we also factor in drivers’ attention level and its effect on their mode preferences. Additionally, we do not restrict ourselves to visual and auditory modes and include other modes such as tactile modes and different types of visual and auditory modes that are omitted from their work. However, unlike Du et al. (2021), we do not consider or test combinations of different modes of communication.
2.2. Adaptive Human-Automation Interaction
@@
Human-automation adaption is a major field of study at the intersection Psychology and Automotive Sciences. Adaptive systems could benefit human performance if the interaction between the human-automation team was adapted to their task and contextual demands (Hancock et al., 2013). These adaptive systems would decrease the cognitive load on drivers leading to an improved driving experience (Gomaa et al., 2022). Such adaptive systems would also vastly improve driver engagement as humans tend to lose interest in the task at hand if automation does not adapt quickly (Yi et al., 2020). Other works have modeled human-automation interaction with the use of Hidden Markov Models and Markov Decision Processes (Janssen et al., 2019; Jokinen and Kujala, 2021; Hasan et al., 2023). These models are trained using simulated data but are yet to be tested with human drivers.
@@
Personalizing ADAS has been popular and proposed to improve driving experiences (Hasenjäger et al., 2020; Boverie et al., 2011). In particular, Boverie et al. (2011) propose DrivEasy, a framework for adaptive ADAS that behaves similarly to AdaCoM systems. While DrivEasy provides an overview for such an adaptive ADAS framework, no information about the design of adaptive communication is provided. Additionally, driver sentiments towards such a system and its adaptive communication are not highlighted. In our work, we directly address this gap by testing driver comfort and perceived safety with such AdaCoM systems.
@@
Figure 2. An overview of the survey: Participants first fill out a demographics questionnaire, followed by a driving experience questionnaire. Then, participants are randomly assigned one of three tracks that contain distinct videos showcasing the 5 different driving scenarios under different environmental conditions and augmentations. Finally, the survey ends with a questionnaire on their need for and comfort with a system with adaptive communication.
[…]
[A flow chart describing the procedure for the user study.]A flow chart describing the procedure for the user study.
2.3. Automotive Study Methodologies
3. Method
@@
In this section, we describe the procedure and contents of our study. We first present an overview of the user . Then, we explain the rationale behind the design of the videos and our survey questionnaires. This study was approved by the institutional review board at the authors’ institute.
3.1. Procedure Overview
@@
The study11footnotemark: 1 was administered online through Google Forms. Participants in the study followed the shown in Figure 2. To begin, participants were asked to fill an informed consent form. After submitting the form, they were directed to a demographics questionnaire and a driving experience questionnaire [^2] Details regarding the demographics and driving experience questionnaires presented in Appendix A.. Only participants who were above the age of 18, resided in the United States, and held a driver’s license were allowed to proceed further. Participants were then randomly assigned one of three video tracks. Each video track consisted of distinct videos and a questionnaire for the scenario shown in each video. Further details regarding the videos and the questionnaires are presented in Section 3.2. Participants were allowed to view the videos multiple times before answering the questionnaire. Following the video track, participants filled out a questionnaire regarding their opinion on an adaptive communication module for vehicles (discussed in Section 3.3).
3.2. Videos and Modes of Communication Questionnaire
@@
Table 1. The detailed video sequences for each track. We show the driving scenario, weather conditions, and if the video was augmented. The slight dominance of scenarios depicting rain in Track 3 was completely incidental.
@@
Scenario Weather Aug. Scenario Weather Aug. Scenario Weather Aug.
@@ 10 Lane Deviation Rain Lane Deviation Night Pedestrian Collision Rain
, participants were shown videos from a driver’s perspective simulated in the CARLA Driving Simulator (Dosovitskiy et al., 2017). All videos can be viewed on the project webpage11footnotemark: 1. The videos depict realistic, high-risk scenarios, which could not safely be collected in a real vehicle. As seen in Figure 1(c), the layout of the videos was designed to show the front view through the windshield, and both side views and the rear-view through their respective mirrors.
@@
Each video included one of five driving scenarios at varying levels of risk. The driving scenarios, in order of decreasing risk were: (1) Pedestrian Collisions; (2) Car Collisions; (3) Lane Deviation; (4) Blind Spot Obstruction; (5) Stagnant at a Green Light. The five scenarios were selected due to the ubiquity of ADAS developed to address them (e.g. collision detectors, lane keeping assist systems). The weather in the videos varied between: (a) clear day; (b) rainy day; and (c) night. In total, 15 videos11footnotemark: 1 ($3$ weather conditions $\times$ $5$ scenarios) were produced. The number of pedestrians and vehicles and the environment setting (urban center or highways) were also varied across each video i.e. each video is unique with overlaps in only the type of scenario and the weather condition.
@@
Environmental variations and other agents in the simulation were used to introduce realistic distractions. Particularly, multiple video editing techniques extrapolated from prior work focusing on driver inattention were used to simulate distractions such as: (1) drowsiness; (2) hyper-focus on the road; and (3) noticeable agents in the environment (Health, 2023; Shiferaw et al., 2014). These distractions were chosen due to their occurrence in driver distraction literature and the ease of implementing them in a simulated environment. Drowsiness was simulated by continually splicing the video with a plain black screen with an increasing duration and frequency to mimic an increase in blinking rate and the change in perception associated with drowsy driving (Health, 2023). Inspired by existing literature, we used editing techniques such as zooming in, cropping out, and selectively blurring areas on the videos to simulate hyper-focus on areas in the environment (Shiferaw et al., 2014). We also used other dynamic agents in the environment such as children, emergency vehicles, and pedestrians with unique behaviors as sources of distraction outside vehicles. The videos were edited to blur out all views except the aforementioned elements and/or dynamic agents of interest.
@@
Each of the 15 videos were augmented in only one of the styles described above to produce 15 pairs of videos ($3$ weather conditions $\times$ $5$ scenarios $\times$ 2 versions). Therefore, a total of 30 videos were produced and ordered in each of the three tracks (Table 1). The 15 pairs of videos were divided into three tracks such that each driving scenario appeared exactly once in every track with no overlapping videos across tracks. In particular, each track comprised of 5 videos covering each driving scenario and their augmented versions. All augmented versions of videos appeared before the unedited videos. This ordering reduced bias by ensuring that participants did not have a complete view of the scenarios in the augmented versions that they could obtain in the original unedited videos. The videos were also distributed as to have maximum coverage in all tracks for the differing weather conditions and distraction augmentations. Therefore, each scenario was viewed and evaluated independently by every participant.
@@
Table 2. The Options for the Modes of Communication.
@@
As with similar online studies (see Section 2.3), the participants were asked to answer a mode preference questionnaire after carefully watching each video by imagining themselves as the driver of the vehicle. The questionnaire involved ranking the modes of communication listed in Table 2 and indicating if their vehicle should automatically takeover control. The modes in Table 2 were selected due to their popular use in vehicles and prior literature (Campbell et al., 2018). Participants were required to choose only their top three preferred modes for the scenario taking place in every video shown. A Likert scale was used to gather responses for sentiments regarding automatic takeovers for each video.
The questionnaire was designed to obtain the drivers’ mode preferences across real-world scenarios with differing risk and under distractions. To minimize the effects of individual biases, inferences from the videos were extrapolated using a within-subjects strategy. For this portion of the study, the dependent variables were the ranks associated with each mode and the independent variables were attributes associated with each video i.e. the driving scenario, the weather conditions, and if and how the video was augmented.
We believe that these tests are robust due to the considerations in the design of the video tracks. Further details regarding the evaluation are discussed in Section 4.
3.3. Adaptive Communication Module Questionnaire